A 25 meter ladder leans against a vertical wall. The foot of the ladder is 20 meter from the base of the wall. If the foot is moved 13 meter closer to the wall, how far does the top of the ladder move up the wall?
Answer by fractalier(6550)

You can deduce the ladder's original height from the Pythagorean Theorem...
h^2 = 25^2 - 20^2
h^2 = 225
h = 15 feet
If its base is moved 13 feet closer it will be 7 feet from the wall.
The new height is again found the same way...
h^2 = 25^2 - 7^2
h^2 = 576
h = 24 feet
and so moves up 24 - 15 =
9 feet